OpenStudy (surana):

Given a = 2, b = -2 and k = 3, what is the equation of the graph if the parent function is y= /-x?

OpenStudy (surana):

Technically, it's \[y=\sqrt{x}\]

OpenStudy (surana):

@bohotness , @misty1212

OpenStudy (bohotness):

do you know how to do tha/T

OpenStudy (surana):

Sort of, I need help with things like this.

OpenStudy (bohotness):

okay

OpenStudy (surana):

Yes.

geerky42 (geerky42):

Exactly what do a, b, and k stand for?

geerky42 (geerky42):

Is it some sort of transformations?

OpenStudy (surana):

a, b and k stand for, respectively, 2, -2 and 3.

geerky42 (geerky42):

I know, but what do they represent?

geerky42 (geerky42):

let me guess, a is x-shift and b is y-shift, and k is vertical scale? As in \(k\sqrt{-(x-a)}+b\)?

OpenStudy (surana):

Variables for the equation.

OpenStudy (surana):

That might be it, actually.

geerky42 (geerky42):

Ok, then just plug in value and you will have your equation, I guess

geerky42 (geerky42):

I made assumption, so I could be wrong.

OpenStudy (surana):

So I put the numbers into the places where the variables themselves are and that's my answer?

geerky42 (geerky42):

I think so. your question doesn't give enough information for us

OpenStudy (surana):

Well, that is the entirety of the question.

OpenStudy (surana):

3\[3\sqrt{-x(-2)} + (-2)\]

OpenStudy (surana):

So it's basically that?

geerky42 (geerky42):

\(3\sqrt{(-x-2)}-2\) but this is only if general function is \(k\sqrt{(-x-a)}+b\)

geerky42 (geerky42):

It depend on where a, b, and k is in general equation. I think you should be given that in maybe previous questions or something?

OpenStudy (surana):

I think it follows the general equation.

geerky42 (geerky42):

well, whether it is, just plug in and simplify if needed, and you are good.

geerky42 (geerky42):

couldn't be much help, sorry

OpenStudy (surana):

It's alright. thank you.